Matthews & Bros. is a local landscape construction company. In analyzing financial performance, the cost accountant compares actual results with a flexible budget. The standard direct labor rates used in the flexible budget are established each year at the time the annual plan is formulated and held constant for the entire year.

 

The standard direct labor rates in effect for the current fiscal year and the standard hours allowed for the actual output of work for July are shown in the following schedule.

 

Worker Classification
Standard Direct Labor Rate per Hour
Standard Direct Labor-Hours Allowed for Output
I
$44

1,710
II

39

2,280
III

29

2,850

 

The wage rates for each labor class increased under the terms of a new contract. The standard wage rates were not revised to reflect the new contract.

 

The actual direct labor-hours worked and the actual direct labor rates per hour experienced for the month of July were as follows.

 

Worker Classification
Actual Direct Labor Rate per Hour
Actual Direct Labor-Hours
I
$48

1,800
II

42

2,400
III

30

2,950

 

Required:

Calculate the dollar amount of the total direct labor variance for July for Matthews & Bros. and break down the total variance into the following components:

 

a. Direct labor price and efficiency variances.

b. Direct labor mix and yield variances.
Answer #1

step 1) 1,710+ 2,280+ 2,850= 6,840 total Standard Direct Labor-Hours Allowed for Output

step 2) 1,710/ 6,840= 0.25

step 3) 2,280/ 6,840= 0.3333333333

step 4) 2,850/ 6,840= 0.4166666667 or 0.417 

step 5) 

Actual: ($48 × 1,800) + ($42 × 2,400) + ($30 × 2,950) = $275,700


Flexible Production Budget: ($44 × 1,710) + ($39 × 2,280) + ($29 × 2,850) = $246,810


(SP × AQ): ($44 × 1,800) + ($39 × 2,400) + ($29 × 2,950) = $258,350


(SP × ASQ): ($44 × 0.250a × 7,150) + ($39 × 0.333a × 7,150) + ($29 × 0.417a × 7,150) = $257,996


Part A answer

Direct labor price variance = $275,700 – $258,350 = $17,350 U

Efficiency variances = $258,350 – $246,810 = $11,540 U


Part B answer

Yield variance = $257,996 – $246,810 = $11,186 U

Direct labor mix = $258,350 – $257,996 = $354 U
Answer #2

ignore the a's in (SP × ASQ) as they are ahowing that they are from the results of steps 2-4